OpenStudy (anonymous):

The taylor series for f(X)=sin(X) at a=pie/3 is sigma (cn(x-pie/3)^(n)). find c0,c1,c2,c3,c4

OpenStudy (experimentx):

what is the second statement??

OpenStudy (anonymous):

oh just find each C for n=0,1,2,3,4

OpenStudy (anonymous):

lol "pie"

OpenStudy (experimentx):

\(sin(x) = sin(a) + cos(a) x - sin(a) x^2/2 - cos(a)/x^3/3! + sin(a) x^4/4 + ...\) \(sin(x) = sin(\pi/3) + cos(\pi/3) x - sin(\pi/3) x^2/2 - cos(\pi/3)/x^3/3! + sin(\pi/3) x^4/4! + ...\) From above, it can be found by comparing

OpenStudy (experimentx):

\( c0 = \sqrt{3}/2\), \(c1 = 1/2\), .. ..

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{2}+\frac{1}{2} \left(x-\frac{\pi }{3}\right)-\frac{1}{4} \sqrt{3} \left(x-\frac{\pi }{3}\right)^2-\frac{1}{12} \left(x-\frac{\pi }{3}\right)^3+\frac{\left(x -\frac{\pi }{3}\right)^4}{16 \sqrt{3}}+\frac{1}{240} \left(x-\frac{\pi }{3}\right)^5+O\left(\left( x-\frac{\pi }{3}\right)^6\right) \]

OpenStudy (experimentx):

wow?? how did you do that??

OpenStudy (experimentx):

So, i must have been (x-pi/3) instead of x how will you get the coefficients of x's ... isn't it going to infinity??